# Memory-Augmented Generative AI for Real-time Wireless Prediction in Dynamic Industrial Environments

### v1

@@ 3.4 Joint Loss Calculation and Optimization
The entire architecture is optimized end-to-end using a single joint loss function. This unified optimization strategy ensures synergistic learning, compelling both the generative capabilities of the Dynamic_VAE and the predictive accuracy of the ConvLSTM to evolve in concert towards the final SINR prediction task. The combined loss function $L_{Total}$ comprises three distinct terms:
  1. 1.
@@
    VAE Reconstruction Loss ($L_{VAE\_recon}$): This term quantifies the accuracy of the VAE’s reconstructed SINR maps against the true SINR sequence. It is calculated as the Mean Squared Error (MSE) between the VAE’s predicted sequence (vae_pred_seqvae\_pred\_seq) and the ground truth SINR sequence ($y_{seq}$), both typically of shape [Batch, Sequence_Len, 1, H, W]:
    $\begin{split}L_{\text{VAE\_recon}}=&\frac{1}{N\cdot T\cdot H\cdot W}\\ &\cdot\sum_{b,t,h,w}(\text{vae\_pred\_seq}_{b,t,h,w}-y_{\text{seq},b,t,h,w})^{2}\end{split}$ (14)
@@
    where $N$ is the batch size, $T$ is the sequence length, $H$ is height, and $W$ is width.
  2. 2.
@@
    VAE KL Divergence Loss ($L_{KLD}$): This regularization term ensures that the latent distributions generated by the VAE’s encoder ($\mathcal{N}(\mu_{t},\sigma_{t}^{2})$) are pushed towards a standard normal distribution ($\mathcal{N}(0,I)$). This promotes a smooth, continuous, and disentangled latent representation space, facilitating better generalization and generation. For a multivariate Gaussian distribution, the Kullback-Leibler Divergence (KLD) from $\mathcal{N}(\mu,\Sigma)$ to $\mathcal{N}(0,I)$ is given by:
    $L_{KLD}=\frac{1}{2}\sum_{t=1}^{\text{Sequence\_Len}}\sum_{i=1}^{D}(\sigma_{t,i}^{2}+\mu_{t,i}^{2}-1-\log(\sigma_{t,i}^{2}))$ (15)
@@
    where $D$ is the dimension of the latent vector $z_{t}$. This term effectively penalizes deviations of the latent distribution from the prior.
  3. 3.
@@
    LSTM Prediction Loss ($L_{LSTM\_pred}$): This term directly measures the accuracy of the ConvLSTM’s final spatio-temporal predictions. It is calculated as the MSE between the ConvLSTM’s output sequence (outputs_seqoutputs\_seq) and the true SINR sequence ($y_{seq}$):
    $\begin{split}L_{\text{LSTM\_pred}}=&\frac{1}{N\cdot T\cdot H\cdot W}\\ &\cdot\sum_{b,t,h,w}(\text{outputs\_seq}_{b,t,h,w}-y_{\text{seq},b,t,h,w})^{2}\end{split}$ (16)
The total loss function $L_{Total}$ is a weighted sum of these three components:
@@
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters that control the relative importance of each loss term during training. These weights are typically determined through empirical tuning (e.g., grid search or Bayesian optimization) to achieve optimal model performance. This unified optimization ensures that the entire architecture is trained collaboratively, leveraging the strengths of both generative modeling and sequence prediction for the ultimate task of precise spatio-temporal SINR prediction.
4 Experimental Evaluation
@@ 4.1 Experiment Methodology
Following the chronological split, we transformed the datasets into time-series sequences using a sliding window approach. For a given ‘sequence_length‘, each input sample for the model consisted of a series of multi-channel images (Euclidean distance, Permeability, and AP location maps) corresponding to that time window. The corresponding label for each sequence was the series of SINR heatmaps for the same time steps. The window was slid by one time step at a time, generating overlapping sequences for training and evaluation, which allowed the model to learn the temporal dynamics between consecutive frames (see Figure 2).
@@ 4.2 Implementation Details
4.3 Ablation Studies: Quantifying the Contribution of the LMM and the Hybrid Architecture
@@
Figure 4: t-SNE of WISVA Latent Space [29].
The comparison between the t-SNE visualization of the Evo-WISVA latent space (Figure 3) and the baseline WISVA latent space (Figure 4) provides crucial insight into the structural benefits introduced by the Attention-driven Latent Memory Module (LMM) and the ConvLSTM components of Evo-WISVA.
  1. 1.
    Latent Space Scale and Separation (Architectural Implication): A primary distinction is the scale: Evo-WISVA spans approximately $\mathbf{\pm 150}$ on both axes, whereas WISVA spans only $\mathbf{\pm 15}$.
    • WISVA (Small Scale): As a standard VAE, its latent vectors ($\mathbf{z}$) are heavily constrained by the KL divergence loss, forcing adherence to a compact Standard Gaussian prior ($\mathcal{N}(0,\mathbf{I})$). This regularization results in a denser, smaller-magnitude latent manifold.
@@
    • Evo-WISVA (Large Scale): The complex, memory-augmented architecture generates latent means ($\boldsymbol{\mu}$) that are widely separated and less constrained by a strict Gaussian prior. The LMM and attention mechanism enforce significant distances between distinct contextual states, translating into a substantially larger scale in the t-SNE projection. This separation confirms the model’s success in robustly encoding highly distinct environmental modes.
  2. 2.
    SINR Organization and Disentanglement:
    • Evo-WISVA: Exhibits tight, highly homogeneous clusters that organize clearly and monotonically along the t-SNE dim 2 (y-axis). High SINR ($\approx 25$ dB) is concentrated at the top, and low SINR ($\approx 0$ dB) at the bottom. This demonstrates the model’s precise feature differentiation even across dynamic dataset.
    • WISVA: Presents a diffuse, scattered distribution with substantial cluster overlap. The SINR coloring shows a wider range but lacks structural organization: high, medium, and low SINR points are heavily mixed, indicating a lower degree of SINR feature isolation.
  3. 3.
@@
    Conclusion on Latent Quality: The Evo-WISVA latent space is demonstrably better organized. The distinct clustering and clear ordering of the SINR metric along a single dimension signify that the LMM and hybrid design successfully disentangle the SINR feature from other encoded factors. This highly structured, high-separation encoding is critical for the subsequent ConvLSTM to perform the accurate, context-aware temporal forecasting required to meet Ultra-Reliable Low-Latency Communication (URLLC) requirements.
4.4 Evo-WISVA: Comparison with State-of-the-Art Models
@@
To rigorously evaluate the efficacy of the proposed Evo-WISVA architecture, its performance was benchmarked against several state-of-the-art deep learning models for spatiotemporal wireless channel prediction. The selected baselines include a standalone ConvLSTM model [33], a Wireless Infrastructure for Smart Warehouse using Variational Autoencoder (WISVA) model [29], an LSTM-based Link Quality Indicator (LSTM-LQI) model [34], and a GRU-LSTM model [35]. This comparative analysis is designed to highlight the advantages conferred by Evo-WISVA’s synergistic and memory-augmented design.
@@
Figure 5 illustrates the average reconstruction error per pixel (in dB) for each model as a function of the input sequence length. The data unequivocally demonstrate that the Evo-WISVA model consistently achieves a significantly lower prediction error across all tested sequence lengths. For instance, at a sequence length of 8, Evo-WISVA’s average reconstruction error is 47.6% lower than that of the second-best model (ConvLSTM) and 78.4% lower than the LSTM-LQI model.
@@
The superior performance of Evo-WISVA is primarily attributed to its innovative hybrid architecture. The memory-augmented VAE component provides a robust and context-aware initial estimate of the SINR map by learning a compressed, latent representation of the input data. The subsequent ConvLSTM module leverages this enriched spatiotemporal information to perform a precise refinement of the predicted SINR map. This two-stage process effectively mitigates the inherent noise and stochasticity of the wireless channel, which simpler, single-stage models are less equipped to handle.
By effectively decoupling the data compression and spatiotemporal prediction tasks, the Evo-WISVA model achieves a higher degree of predictive accuracy, positioning it as a highly effective solution for complex wireless channel forecasting applications.
4.5 Performance on Limited Time-Series Data
@@
Figure 6 and Figure 7 illustrate the model’s Mean Absolute Error (MAE) and Root Mean Square Error (RMSE), respectively, for both the full validation dataset and the limited time-series subset. The results demonstrate that the model maintains a consistent level of performance, with only minor fluctuations, confirming its robust generalization capability.
Figure 6: MAE comparison of Evo-WISVA on the overall test dataset versus a limited time-series dataset.
Figure 7: RMSE comparison of Evo-WISVA on the overall test dataset versus a limited time-series dataset.
The model’s resilience to data volume variations is crucial for a digital twin, as it must maintain predictive accuracy when operating on-the-fly with continuous data streams from a live network environment. The results confirm that the model’s performance is not contingent on a large, diverse data sample at runtime, which is a key requirement for practical deployment.
4.6 Real-time Prediction Performance and Scalability Analysis
@@
4.7 Generalization to Increased Environmental Complexity
One of the most significant challenges in developing a wireless digital twin for a smart warehouse is handling the unpredictability introduced by dynamic environmental elements. To assess the robustness and generalization ability of the proposed Evo-WISVA model, we designed a transfer evaluation setting. The model was trained exclusively on a dataset containing a single moving shelf and subsequently tested on unseen datasets where 2, 4, 6, and 10 shelves were moving simultaneously. This evaluation is particularly important, as the number of moving shelves directly correlates with the degree of multipath interference and temporal variability in the wireless channel.
@@ 4.7 Generalization to Increased Environmental Complexity
The memory-augmented VAE component proved essential for disentangling the stable physical properties of the warehouse layout from the stochastic, time-varying perturbations induced by moving shelves. Coupled with the ConvLSTM’s ability to model sequential dependencies, the architecture was able to preserve high accuracy under conditions that were far more complex than those encountered during training. This level of robustness is crucial for the deployment of predictive digital twins in real-world Industry 4.0 environments, where environmental dynamics cannot be exhaustively pre-encoded in training datasets.
5 Conclusion
In this paper, we introduced Evo-WISVA, a novel synergistic deep learning architecture for a Memory-Augmented Real-Time 2D Wireless Digital Twin, specifically designed for dynamic SINR prediction in Industry 4.0 warehouse environments. By integrating a memory-augmented Variational Autoencoder (VAE) with a ConvLSTM in an end-to-end jointly trained framework, our solution effectively addresses the challenges posed by mobile obstacles and complex RF propagation. We demonstrated that Evo-WISVA can generate accurate spatio-temporal SINR predictions while achieving significant computational acceleration, establishing it as a foundational technology for proactive wireless resource management and predictive digital twins.
Industrial environments are inherently dynamic, rendering static network planning insufficient. The proposed model enables real-time foresight into wireless performance, allowing network operators to transition from reactive troubleshooting to proactive resource management. This represents a paradigm shift in industrial connectivity, as Evo-WISVA supports decision-making for dynamic beamforming, intelligent routing, and other adaptive network operations. By emulating complex wireless effects quickly and accurately, it reduces reliance on time-consuming physical simulations or measurements, further enhancing operational efficiency and reliability.
Future research directions include extending the model to multi-AP scenarios with dynamic handovers and interference management, integrating real-time sensor data (e.g., LiDAR-based obstacle detection) to improve prediction fidelity, and deploying the architecture on resource-constrained edge devices such as industrial gateways or autonomous mobile robots. Additionally, exploring the transferability of Evo-WISVA to other dynamic environments beyond warehouses will be key to establishing a generalizable framework for predictive wireless digital twins.

### v3

@@ III-D Joint Loss Calculation and Optimization
The entire architecture is optimized end-to-end using a single joint loss function. This unified strategy ensures synergistic learning, compelling both the generative capabilities of the Dynamic_VAE and the predictive accuracy of the ConvLSTM to evolve in concert. The total objective $L_{Total}$ comprises three distinct terms:
VAE Reconstruction Loss ($L_{VAE\_recon}$)
@@
This term quantifies the accuracy of the VAE’s reconstructed SINR maps against the ground truth. It is calculated as the Mean Squared Error (MSE) between the predicted sequence (vae_pred_seqvae\_pred\_seq) and the true sequence ($y_{seq}$):
$L_{\text{VAE\_recon}}=\frac{\sum_{b,t,h,w}(\text{vae\_pred\_seq}_{b,t,h,w}-y_{\text{seq},b,t,h,w})^{2}}{N\cdot T\cdot H\cdot W}$ (14)
@@
where $N$, $T$, $H$, and $W$ represent batch size, sequence length, height, and width, respectively.
VAE KL Divergence Loss ($L_{KLD}$)
@@
This regularization term ensures that the latent distributions $\mathcal{N}(\mu_{t},\sigma_{t}^{2})$ are pushed towards a standard normal prior $\mathcal{N}(0,I)$. This promotes a smooth, continuous, and disentangled latent space. For a multivariate Gaussian distribution, the KLD is formulated as:
$L_{KLD}=\frac{1}{2}\sum_{t=1}^{T}\sum_{i=1}^{D}(\sigma_{t,i}^{2}+\mu_{t,i}^{2}-1-\log(\sigma_{t,i}^{2}))$ (15)
@@
where $D$ is the latent dimension. This term penalizes deviations from the prior, preventing posterior collapse.
LSTM Prediction Loss ($L_{LSTM\_pred}$)
@@
This term measures the final spatio-temporal predictive accuracy of the ConvLSTM. It is the MSE between the model outputs (outputs_seqoutputs\_seq) and the true SINR maps ($y_{seq}$):
$L_{\text{LSTM\_pred}}=\frac{\sum_{b,t,h,w}(\text{outputs\_seq}_{b,t,h,w}-y_{\text{seq},b,t,h,w})^{2}}{N\cdot T\cdot H\cdot W}$ (16)
The total loss $L_{Total}$ is a weighted sum of these components, where $\alpha$, $\beta$, and $\gamma$ are hyperparameters determined via empirical tuning to balance generative stability with predictive precision:
@@
IV Experimental Evaluation
@@ IV-A Experiment Methodology
Dataset Credibility and Validation: The use of ns-3 for 60 GHz industrial modeling is well-supported by recent comparative studies. Specifically, the authors [34] cross-validated the ns-3 802.11ad/ay modules against a physical 60 GHz testbed, reporting an average error of only 2.43% and a standard deviation of 1.33 dBm. Furthermore, our environment utilizes the Quasi-Deterministic (Q-D) channel model, which authors [35] demonstrated to be mathematically equivalent to deterministic ray-tracing in terms of 3D obstacle specification and LoS/NLoS accuracy. This ensures that our training data captures site-specific multipath and shadowing effects with high physical fidelity.
Following the chronological split, we transformed the datasets into time-series sequences using a sliding window approach. For a given ‘sequence_length‘, each input sample for the model consisted of a series of multi-channel images (Euclidean distance, Permeability, and AP location maps) corresponding to that time window. The corresponding label for each sequence was the series of SINR heatmaps for the same time steps. The window was slid by one time step at a time, generating overlapping sequences for training and evaluation, which allowed the model to learn the temporal dynamics between consecutive frames (Figure 2).
@@ IV-B Implementation Details
IV-C Ablation Studies: Quantifying the Contribution of the LMM and the Hybrid Architecture
@@
Figure 4: t-SNE of WISVA Latent Space.
@@
The comparison between the t-SNE visualizations of the Evo-WISVA latent space (Fig. 3) and the baseline WISVA manifold (Fig. 4) provides crucial insight into the structural benefits of the Attention-driven Latent Memory Module (LMM) and ConvLSTM components. Latent Space Scale and Separation: A primary distinction is observed in the manifold scale; Evo-WISVA spans approximately $\pm 150$ on both axes, whereas WISVA spans only $\pm 15$. As a standard VAE, WISVA’s latent vectors $\mathbf{z}$ are heavily constrained by the KL divergence loss, forcing adherence to a compact standard Gaussian prior $\mathcal{N}(0,\mathbf{I})$ and resulting in a denser, smaller-magnitude manifold. Conversely, the complex, memory-augmented Evo-WISVA architecture generates latent means $\boldsymbol{\mu}$ that are widely separated and less constrained by the strict Gaussian prior. The LMM and attention mechanism enforce significant distances between distinct contextual states, confirming the model’s success in robustly encoding highly distinct environmental modes.
@@
SINR Organization and Disentanglement: Regarding feature organization, Evo-WISVA exhibits tight, highly homogeneous clusters that align monotonically along the second t-SNE dimension, with high SINR ($\approx 25$ dB) and low SINR ($\approx 0$ dB) concentrated at opposite poles. This precise differentiation contrasts sharply with WISVA, which presents a diffuse, scattered distribution with substantial cluster overlap and poor structural organization. The distinct clustering and clear ordering of the SINR metric in Evo-WISVA signify that the LMM and hybrid design successfully disentangle the SINR feature from other encoded environmental factors. This highly structured encoding is critical for the subsequent ConvLSTM to perform the accurate, context-aware temporal forecasting required for URLLC requirements.
Conclusion on Latent Quality: These findings empirically validate Lemma 1, indicating that Evo-WISVA successfully expands the spectral representation of the dynamic SINR manifold. By filtering dominant static variance through the VAE and short-circuiting the temporal gradient via the LMM, the model achieves the disentanglement visible in Fig. 3, where SINR levels are linearly separable along a single axis. This separation proves that the model has avoided the common pitfall of posterior collapse, preserving high-frequency spatial features necessary for industrial forecasting.
TABLE I: Sensitivity Analysis of Joint Loss Hyperparameters on Prediction Accuracy.
Case $\alpha$ (VAE) $\beta$ (KLD) $\gamma$ (LSTM) MAE (dB)
1 1.0 0.001 1.0 1.7451
2 1.0 0.01 1.0 2.9091
3 1.0 0.1 1.0 2.0498
4 0.5 0.01 1.0 2.1149
5 1.0 0.01 0.5 2.6189
6 1.5 0.01 1.0 1.9485
7 1.0 1.0 1.0 2.0590
The sensitivity analysis, summarized in Table I, reveals that the Evo-WISVA framework is highly sensitive to the variational bottleneck weight $\beta$. Specifically, Case 1 ($\beta=0.001$) achieved the lowest error (MAE), suggesting that excessive KL-regularization in dynamic industrial environments can lead to “posterior collapse,” where subtle RF shadowing details are smoothed over. Furthermore, the performance gain from Case 4 to Case 6 demonstrates that the spatial reconstruction weight $\alpha$ is a primary determinant of accuracy, validating the VAE’s role as a robust spatial prior. The ConvLSTM ($\gamma$) contributes to temporal refinement, though its efficacy is contingent upon the stability of the latent features provided by the VAE.
IV-D Evo-WISVA: Comparison with State-of-the-Art Models
@@
To rigorously evaluate the efficacy of the proposed Evo-WISVA architecture, its performance was benchmarked against several state-of-the-art deep learning models for spatiotemporal wireless channel prediction. The selected baselines include a standalone ConvLSTM model [36], a Wireless Infrastructure for Smart Warehouse using Variational Autoencoder (WISVA) model, an LSTM-based Link Quality Indicator (LSTM-LQI) model [37], and a GRU-LSTM model [38]. Furthermore, to provide a broader context of architectural evolution, Table II provides a comparative framework of these and other stochastic recurrent architectures, highlighting differences in memory design, spatial resolution, and physics retention. This comprehensive comparison is designed to highlight the advantages conferred by Evo-WISVA’s synergistic and memory-augmented design.
@@
Figure 5 illustrates the average reconstruction error per pixel (in dB) for each model as a function of the input sequence length. The data unequivocally demonstrate that the Evo-WISVA model consistently achieves a significantly lower prediction error across all tested sequence lengths.
@@
As synthesized in Table II, the superior performance of Evo-WISVA is primarily attributed to its innovative hybrid architecture. While traditional models like SVG-LP [39] or VRNN [40] often suffer from hidden state drift or limited spatial resolution, the memory-augmented VAE component of Evo-WISVA provides a robust and context-aware initial estimate of the SINR map.
The subsequent ConvLSTM module leverages this enriched spatiotemporal information to perform a precise refinement. This two-stage process—summarized by the “Triple-Stream” input logic in Table II—effectively mitigates the inherent noise and stochasticity of the wireless channel, which simpler, single-stage models are less equipped to handle.
TABLE II: Comparative Framework of Stochastic Recurrent Architectures for SINR Sequence Prediction
Model Memory Architecture Spatial Resolution Input Streams Temporal Dynamics Physics Retention
SVG-LP [39] Hidden State only Latent-only Single Encoder Learned Prior Low (Drift)
VRNN [40] Recurrent ($h_{t}$) Latent + RNN Coupled Encoder Conditional Prior Moderate
LSTM-LQI [37] Dual LSTM (Det.+Stoch.) Wavelet-Denoised SNR Det.+Var. Streams CI Boundary Low (Point)
Seq2Seq [38] Enc-Dec (LSTM/GRU) Sequence RSSI Single Signal Recursive Unguided Moderate
ConvLSTM [36] Recurrent ($C_{t},H_{t}$) Conv. Grid Single (Radar/SINR) Causal Conv. Moderate
WISVA [26] VAE Latent Space Grid (152x152) Multi-Physics Static Recon. High
TimeSformer [41] Global Attention Patch-based (ViT) Frame Patches Divided Space-Time High (Global)
PredRNN++ [42] Dual (C+M) + GHU Multi-scale Conv Cascaded Causal Dual-Path Recur. High (Long)
Evo-WISVA Evol. Recur.+Attn. Multi-Res ConvLSTM Triple-Stream Stoch. Attn. Prior Ultra-High
IV-E Visual Analysis of Spatio-Temporal Prediction Quality
To qualitatively assess the Evo-WISVA architecture’s ability to accurately capture the spatial distribution and temporal dynamics of the wireless channel, we present a visual comparison of the predicted SINR heatmap against the ground truth. Figure 6 showcases a randomly selected unseen frame from the test dataset, highlighting the model’s predictive fidelity.
(a) Ground Truth SINR Map
(b) Evo-WISVA Predicted SINR Map
(c) Reconstruction Error Map
Figure 6: Visual comparison of Evo-WISVA’s prediction at an unseen time step. Subfigure (a) is the input ground truth, (b) is the corresponding prediction, and (c) shows the pixel-wise reconstruction error, demonstrating high predictive fidelity across the dynamic environment.
A closer look at the visualizations (Figure 6) directly supports the quantitative metrics. The predicted SINR field demonstrates strong spatial fidelity by accurately tracking the ground truth, effectively anticipating complex environmental effects like shadowing and multipath induced by dynamic obstacles. This accuracy is quantified by the reconstruction error map, which is overwhelmingly characterized by low errors (blue regions, near 0 dB), thus confirming the low Mean Absolute Error (MAE) established previously. Errors that are marginally higher (red regions) are localized to sharp boundaries and areas undergoing rapid line-of-sight (LOS) to non-line-of-sight (NLOS) transitions, which represent the most challenging predictive regions. In totality, this qualitative assessment reinforces the robustness and precision conferred by the novel memory-augmented, hybrid Evo-WISVA architecture.
IV-F Performance on Limited Time-Series Data
@@
Figure 7 illustrate the model’s Mean Absolute Error (MAE) and Root Mean Square Error (RMSE), for both the full validation dataset and the limited time-series subset. The results demonstrate that the model maintains a consistent level of performance, with only minor fluctuations, confirming its robust generalization capability.
(a) MAE Comparison
(b) RMSE Comparison
Figure 7: Predictive performance of Evo-WISVA on the overall test dataset versus a limited 100-sample time-series dataset.
The model’s resilience to data volume variations is crucial for a proactive network management, as it must maintain predictive accuracy when operating on-the-fly with continuous data streams from a live network environment. The results confirm that the model’s performance is not contingent on a large, diverse data sample at runtime, which is a key requirement for practical deployment.
IV-G Real-time Prediction Performance and Scalability Analysis
@@
IV-H Generalization to Increased Environmental Complexity
One of the most significant challenges in developing a wireless prediction system for a smart warehouse is handling the unpredictability introduced by dynamic environmental elements. To assess the robustness and generalization ability of the proposed Evo-WISVA model, we designed a transfer evaluation setting. The model was trained exclusively on a dataset containing a single moving shelf and subsequently tested on unseen datasets where 2, 4, 6, and 10 shelves were moving simultaneously. This evaluation is particularly important, as the number of moving shelves directly correlates with the degree of multipath interference and temporal variability in the wireless channel.
@@ IV-H Generalization to Increased Environmental Complexity
The memory-augmented VAE component proved essential for disentangling the stable physical properties of the warehouse layout from the stochastic, time-varying perturbations induced by moving shelves. Coupled with the ConvLSTM’s ability to model sequential dependencies, the architecture was able to preserve high accuracy under conditions that were far more complex than those encountered during training. This level of robustness is crucial for the deployment of predictive systems in real-world Industry 4.0 environments, where environmental dynamics cannot be exhaustively pre-encoded in training datasets.
V Conclusion
In this paper, we introduced Evo-WISVA, a novel synergistic deep learning architecture for a Memory-Augmented Wireless Predictive System, specifically designed for dynamic SINR forecasting in Industry 4.0 warehouse environments. By integrating a memory-augmented Variational Autoencoder (VAE) with a ConvLSTM in an end-to-end jointly trained framework, our solution effectively addresses the challenges posed by mobile obstacles and complex RF propagation. We demonstrated that Evo-WISVA can generate accurate spatio-temporal SINR predictions while achieving significant computational acceleration, establishing it as a foundational technology for proactive wireless resource management and advanced predictive systems.
Industrial environments are inherently dynamic, rendering static network planning insufficient. The proposed model enables real-time foresight into wireless performance, allowing network operators to transition from reactive troubleshooting to proactive resource management. This represents a paradigm shift in industrial connectivity, as Evo-WISVA supports decision-making for dynamic beamforming, intelligent routing, and other adaptive network operations. By emulating complex wireless effects quickly and accurately as a neural surrogate, it reduces reliance on time-consuming physical simulations or measurements, further enhancing operational efficiency and reliability.
Future research directions include extending the model to multi-AP scenarios with dynamic handovers and interference management, integrating real-time sensor data (e.g., LiDAR-based obstacle detection) to improve prediction fidelity, and deploying the architecture on resource-constrained edge devices such as industrial gateways or autonomous mobile robots. Additionally, exploring the transferability of Evo-WISVA to other dynamic environments beyond warehouses will be key to establishing a generalizable framework for predictive industrial informatics.